OpenStudy (anonymous):

Prove algebraically that the equation is an identity. tan(x)+(1/(tan(x)))=1/(sin(x)cos(x))

OpenStudy (anonymous):

\[\tan(\theta)+\frac{1}{\tan \theta}=\frac{1}{\sin(\theta)\cos(\theta)}\] \[\tan(t)+\frac{1}{\tan(t)}=\frac{1}{\sin(t)\cos(t)}\] they are both the same equation.

OpenStudy (anonymous):

tan x + 1/tanx = sinx/cosx + cosx/sinx = sin^2 x + cos^2 x / sinxcosx = 1/sinxcosx

OpenStudy (anonymous):

Alright, let's see if I still have my magic touch. Are you ready to work on this with me?

OpenStudy (anonymous):

yep still here.

OpenStudy (anonymous):

Awesome. First things first: Do you know how else we can show tan? In terms of sin and cosine?

OpenStudy (anonymous):

yeah, there are couple of thing. but they are long equations to write out. him1618 I kind of get it how you got there, but little confused.

OpenStudy (anonymous):

Well, there's one universally simple one where you divide something by something to get tan. Do you remember it? You've gotta memorize this one, it's very important.

OpenStudy (anonymous):

sinx/cosx i believe.

OpenStudy (anonymous):

Perfect! In case you ever forget, I came up with a cheesy way to never forget: There's an island called "sinkland". Here, the people are called "sinkos", or "sincos". And if you meet a sinco, he will be very TAN. Regardless, notice what happens when we switch out tan with sin/cos:

OpenStudy (anonymous):

sin/cos+(1/((sin/cos))=1/(sin*cos)

OpenStudy (anonymous):

Technically you should have x's next to each sin or cos, but I've got a weird thing where I avoid putting them in until I'm all done. It's less messy.

OpenStudy (anonymous):

true! i'm getting that habit also.

OpenStudy (anonymous):

Does it look good so far?

OpenStudy (anonymous):

\[\frac{\sin}{\cos+\frac{1}{\frac{\sin}{\cos}}}=\frac{1}{\sin*\cos}\] does it look like this?

OpenStudy (anonymous):

Actually there should be two fractions on the left side of the equals sign. Can you nudge the (1/sin/cos) to the right of the first fraction?

OpenStudy (anonymous):

don't get it

OpenStudy (anonymous):

Can you explain to me for a moment how to make fractions appear on this site? Never quite figured that part out.

OpenStudy (anonymous):

so it should be (sin/cos)+(1/sin/cos) on the left side?

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

\[\frac{\sin}{\cos}+\frac{\frac{1}{\sin}}{\cos}\]

OpenStudy (anonymous):

perfect.

OpenStudy (anonymous):

Now, one small mathematical "crunch" we will be making. That fraction on the right has two division bars. Do you know the rule about what happens when you do 1/x/y ?

OpenStudy (anonymous):

For example, if you had 1/(1/4), or... 1/.25, what would you do?

OpenStudy (anonymous):

If it's too tricky, lemme know and I'll show you the easy trick for this.

OpenStudy (anonymous):

don't get your question

OpenStudy (anonymous):

i kind of get it. but it looks little confusing.

OpenStudy (anonymous):

Do you seen on the second fraction we have (1/(sin/cos))? We technically should never have a fraction inside a fraction. So we've gotta simplify it. So if I showed you this fraction: 1/(3/5), what would you do?

OpenStudy (anonymous):

If you can show me how to draw fractions here, my examples will make much more sense.

OpenStudy (anonymous):

take the 3/5 only. and add +1 i donk know just guessing.

OpenStudy (anonymous):

No, I'm afraid not.

OpenStudy (anonymous):

i understand your fractions. the complex fractions are fuzzy. where you click "Post" to the left is "Equation" button/option click on it.

OpenStudy (anonymous):

And I just post stuff in the form of 1/4 ?

OpenStudy (anonymous):

\[1/4\]

OpenStudy (anonymous):

Dang it. :/

OpenStudy (anonymous):

Ok, but basically: Imagine 1 divided by (3 divided by 5) Normally, the long way to do it would be to take the 1 and convert it to a fraction: 1/1 and multiply it by the reciprocal of the second fraction: 5/3. The final answer would be 5/3

OpenStudy (anonymous):

Or... as you can tell, if you ever have 1/(some other fraction), the answer is simply "the other fraction flipped upside down"

OpenStudy (anonymous):

ok i get you point. so we ming as well flip that (1/sin/cos) to (cos/sin) which is the same right?

OpenStudy (anonymous):

perfect!

OpenStudy (anonymous):

Which, if you're bored enough to memorize, is also equal to "cotangent" (the inverse of tangent). You'll want to know that for future trig/calculus stuff.

OpenStudy (anonymous):

Alright, so back on topic: we have, sin/cos + cos/sin = 1/cossin

OpenStudy (anonymous):

at some point i'm going to have to know all of them.

OpenStudy (anonymous):

Erm... lemme reword that: sin/cos + cos/sin = 1/cos*sin

OpenStudy (anonymous):

(sin/cos)+(cos/sin)=1/(cos*sin) ???

OpenStudy (anonymous):

perfect. Now, on the left side, we can just cross cancel, right?

OpenStudy (anonymous):

because there's a sin on the top of one and a sin on the bottom of the other, right?

OpenStudy (anonymous):

hit equation on the bottom of your typing screen. and to write a fraction type frac{1}{3} for (1/3)

OpenStudy (anonymous):

Awesome! Thanks.

OpenStudy (anonymous):

\[\frac{\sin}{\cos} + \frac{\cos}{\sin} = \frac{1}{\cos*\sin}\]

OpenStudy (anonymous):

Alright, so can we cancel out the sins and cos to get all ones?

OpenStudy (anonymous):

what are you trying to get me here to? because i'm looking at my question and your explanation and how to make sense of either. it kind of confuses me.

OpenStudy (anonymous):

Ok, that's a good thing then. Because it was a trick question. :P

OpenStudy (anonymous):

and yes with your above equation you cancel them out the get ones

OpenStudy (anonymous):

nononono! There's a plus sign. You can only cross cancel if you're multiplying them together.

OpenStudy (anonymous):

I was using that trick question to make sure you don't accidentally do that on a test. ^_^

OpenStudy (anonymous):

you substitute in one side of the equation for the other side and manipulate it.

OpenStudy (anonymous):

and the other way around.

OpenStudy (anonymous):

No need to do that yet. We're going to make the left side equal 1/(cos*sin) in a few moments.

OpenStudy (anonymous):

Now, how do we add \[\frac{\sin}{\cos} + \frac{\cos}{\sin}\] ?

OpenStudy (anonymous):

beeeehhhiiiivvvveeee

OpenStudy (anonymous):

what do you mean how do we add?

OpenStudy (anonymous):

that equals 1/(sin*cos)

OpenStudy (anonymous):

Eventually it will, yes. But how do we add the fractions?

OpenStudy (anonymous):

As in, what is the rule for adding any two fractions?

OpenStudy (anonymous):

make the denominators alike. so multiply the left denominator by right side. and right denominator by left side of the plus sign

OpenStudy (anonymous):

perfect.

OpenStudy (anonymous):

So what do we multiply \[\frac{\sin}{\cos}\] by to match the fraction next to it?

OpenStudy (anonymous):

Or at least to eventually match the fraction next to it?

OpenStudy (anonymous):

sin

OpenStudy (anonymous):

which would equal in case you are going to ask \[\frac{\sin^2+\cos^2}{\sin*\cos}\]

OpenStudy (anonymous):

yeppers.

OpenStudy (anonymous):

Now, here's where stuff gets tricky if you don't have your identities memorized.

OpenStudy (anonymous):

Do you know of any pythagorean identities?

OpenStudy (anonymous):

hint hint: http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

not memorized. but i've got them written down. by the way is this on problem or extra Info info?

OpenStudy (anonymous):

it's for this problem, I'm afraid.

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

But have no fear! Google will rescue us valiantly.

OpenStudy (anonymous):

http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

Specifically part two, pythagorean identities. Which one applies?

OpenStudy (anonymous):

sin^2+cos^2=1

OpenStudy (anonymous):

awesome. And what do we get?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

BTW, look at what our buddy iman posted on the new questions list. Hee hee.

OpenStudy (anonymous):

Well, I mean after we swap out the identity in our left side problem?

OpenStudy (anonymous):

yeah i've noticed

OpenStudy (anonymous):

And you also noticed what happened to our left hand side equation?

OpenStudy (anonymous):

tan+(1/tan) ???

OpenStudy (anonymous):

Wait. Go back to the last thing we had.

OpenStudy (anonymous):

(sin^2+cos^2) / (sin∗cos)

OpenStudy (anonymous):

\[\frac{\sin^2+\cos^2}{\sin∗\cos}\]

OpenStudy (anonymous):

Remember, you used your identity to swap out your parentheses with a 1, right?

OpenStudy (anonymous):

1/(sin*cos)

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

What is our goal?

OpenStudy (anonymous):

1/(sin*cos)

OpenStudy (anonymous):

Feel accomplished? ;)

OpenStudy (anonymous):

nope. tan+1/tan with what do we substitute that out? i'm confused here.

OpenStudy (anonymous):

No, we did that all already!